128 g of an unknown monatomic gas phase element has just been released into a container of volume 77.26 L at a pressure of 1.0078 atm. If the temperature of the room is 23.0°C, what are the molar mass and the identity of the gas?

What is the density (in g/L) of ethane gas at 743 torr and 14.8°C? The unrounded molar mass of ethane is 30.068 g/mol.
Answer #1

1)

Given:

P = 1.0078 atm

V = 77.26 L

T = 23.0 oC

= (23.0+273) K

= 296 K

find number of moles using:

P * V = n*R*T

1.0078 atm * 77.26 L = n * 0.08206 atm.L/mol.K * 296 K

n = 3.206 mol

mass(solute)= 128 g

use:

number of mol = mass / molar mass

3.206 mol = (128.0 g)/molar mass

molar mass = 39.93 g/mol

This is molar mass of Argon

Answer: 39.93 g/mol, Argon

2)

P= 743.0 torr

= (743.0/760) atm

= 0.9776 atm

T= 14.8 oC

= (14.8+273) K

= 287.8 K

Molar mass of C2H6 = 30.068 g/mol

Lets derive the equation to be used

use:

p*V=n*R*T

p*V=(mass/molar mass)*R*T

p*molar mass=(mass/V)*R*T

p*molar mass=density*R*T

Put Values:

0.9776315789473684 atm * 30.068 g/mol = density * 0.08206 atm.L/mol.K *287.8 K

density = 1.24 g/L

Answer: 1.24 g/L